OpenStudy (anonymous):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!! Eli works at an animal shelter. He needs 40 bags of dog food for the month. There was a buy-4-get-1-free sale on dog food at the grocery shop. Number of bags paid for = 4. 8 , 12 Number of free bags = 1, 2 How many bags of dog food did Eli pay for?

OpenStudy (radar):

If he buys 32 bags, he will get 8 free. That will get him his required 40.

OpenStudy (anonymous):

Thank you!!!!, so much

OpenStudy (anonymous):

he buys 32 bags 8 free right?

OpenStudy (anonymous):

How did you solve this?

OpenStudy (radar):

I am trying to get an equation that would show that lol. 4 = paid for + 1 free one. 4+1=5 40/5 =8 buys (8x4) =32 Gets 8 free one makes 32+8=40

OpenStudy (radar):

I solved it by looking at the progression you started when you asked the question: 4, 8, 12, 16, 20, 24, 28, 32 bought 1, 2, 3, 4, 5, 6, 7, 8 given 5 10 15 20 25 30 35 40 Total